OpenStudy (anonymous):

A wrestling match has eliminations each round. The table below shows the number of wrestlers in each of the first 4 rounds of the tournament: Round (x) 1 2 3 4 Wrestlers f(x) 64 32 18 9 Compute the average rate of change of f(x) from x = 1 to x = 4 and identify the meaning of that rate. A)55; there were 55 fewer wrestlers between rounds 1 and 4 B)3; there were 3 rounds between rounds 1 and 4 C)negative fifty five over three; on average, there were 55 fewer players for every 3 rounds D)negative three over fifty five; on average, there were 3 fewer players for every 55 rounds

OpenStudy (anonymous):

@muscrat123 HEY CAN YOU HELP ME PLEASE

OpenStudy (muscrat123):

@campbell_st

OpenStudy (muscrat123):

WHY DO U KEEP TAGGING ME AFTER I SAY I DONT KNOW HOW TO ANSWER THE QUESTION????

OpenStudy (anonymous):

i sorry i dont see you comment

OpenStudy (muscrat123):

if im not sure, I WILL NOT ANSWER. i dont want to give anybody an incorrect answer

OpenStudy (muscrat123):

i recommend @horsegirl27 @dan815 @jim_thompson5910 @misssunshinexxoxo @campbell_st and @nerlineg

OpenStudy (campbell_st):

average rate of change is the slope of the line use the points (1, 64) and (4, 9)

OpenStudy (anonymous):

yup campbell is right

OpenStudy (anonymous):

yea so what i need to do next?

OpenStudy (campbell_st):

so what did you get as the slope...?

OpenStudy (anonymous):

(1*64)(4*9)= 2304? is right?

OpenStudy (anonymous):

i'm not good at math :(

OpenStudy (campbell_st):

slope = (change in y)/(change in x) sometimes referred to as slope = rise/run

OpenStudy (anonymous):

slope is x2-x1/ y2-y1

OpenStudy (campbell_st):

so the points need to match for x and y slope = (64 - 9)/(1 - 4) what do you get

OpenStudy (anonymous):

-18.3

OpenStudy (anonymous):

?

OpenStudy (campbell_st):

so for your question its avg change = (change in wrestlers)/(change in rounds)

OpenStudy (campbell_st):

well why not leave it as a fraction.... seeing some of the answer choices are fractions

OpenStudy (anonymous):

you have two points plug the x valule with the x vaule and y value with y

OpenStudy (campbell_st):

so as an improper fraction.. what do you thing the avg change of slope would be..?

OpenStudy (anonymous):

i dont know :(

OpenStudy (anonymous):

@campbell_st ???

OpenStudy (anonymous):

are you there?

OpenStudy (anonymous):

please i really need help on this

OpenStudy (campbell_st):

well you had a decimal value avg change = (change in wrestlers)/(change in rounds) = (64 - 9)/(1 - 4)

OpenStudy (anonymous):

=-18.3?

OpenStudy (campbell_st):

can you give the answer as in improper fraction

OpenStudy (anonymous):

how? i dont know how

OpenStudy (campbell_st):

what number is the change in wrestlers what number is the change in rounds...?

OpenStudy (anonymous):

Round (x) 1 2 3 4 Wrestlers f(x) 64 32 18 9

OpenStudy (campbell_st):

change in wrestlers 64 - 9 = avg change = ------------------------------------ -------------- change in rounds 1- 4 = what do you think the missing numbers in the fraction are..?

OpenStudy (anonymous):

-55/ 3

OpenStudy (anonymous):

wrestlers= 55 rounds= -3

OpenStudy (campbell_st):

great so you are looking at choices C or D so read the choices and decide which should be your answer

OpenStudy (anonymous):

-55/-3= 18.3

OpenStudy (anonymous):

D?

OpenStudy (campbell_st):

but your answer choices don't contain a decimal... that's why I asked you to write it as an improper fraction.

OpenStudy (campbell_st):

so you are saying the avg change is \[avg ~change = \frac{-3}{55}\] is that the same as you answer above \[Avg~ change = \frac{-55}{3}\]

OpenStudy (anonymous):

YES

OpenStudy (campbell_st):

well I Think you should choose C the avg change = -55/3

OpenStudy (anonymous):

Yea, thank you so much! can you help me with one more?

OpenStudy (anonymous):

Which of the following options is an equivalent function to f(x) = 3(2)3x? f(x) = 3(8)x f(x) = 24x f(x) = 27(8)x f(x) = 3(8x)

OpenStudy (anonymous):

@Campbell_st

OpenStudy (anonymous):

@nerlineg

OpenStudy (anonymous):

make another post for that answer AND DONT FORGET TO TAG ME :)